As a purchasing manager for Affordable Depot, you collaborate with the marketing, accounting, and engineering departments and strategically manage your procurement activities. You purchase air-conditioner thermostats from its manufacturer @ $45 per unit and sell to customers for $60 per unit. Operating expenses include storage, transportation, marketing, overhead, etc. and add up to $8.00 per unit. According to a reliable forecast, the current annual sales volume is 150,000 thermostats.  You have 3 options: a.Increase the selling price by 2 percent (sales volume will decrease by 2 percent), b.increase sales volume by 2 percent by incurring additional marketing expense of $20,000 (everything else remains the same as the base scenario), or c.work with the manufacturer and get a 3 percent discount on the current cost of goods (everything else remains the same as the base scenario). Compute the percent change in profit for each option with respect to the base profit. Which of the three options will you choose and why? Write your decision in words.
Answer #1

Base profit = sales units * (sale price - purchase cost - expenses) = 150,000 * ($60 - $45 - $8) = $1,050,000

a]

Increased selling price = $60 * (1 + 2%) = $61.20

Decreased sales units = 150,000 * (1 - 2%) = 147,000

Profit = 147000 * ($61.20 - $45 - $8) = $1,205,400

% change in profit = (1205400 - 1050000) / 1050000 = 14.80%

b]

Increased sales volume = 150,000 * (1 + 2%) = 153,000

Profit = 153000 * ($60 - $45 - $8) - $20,000 = $1,051,000

% change in profit = (1051000 - 1050000) / 1050000 = 0.10%

c]
Decreased purchase cost = $45 * (1 - 3%) = $43.65

Profit = 150000 * ($63 - $43.65 - $8) = $1,702,500

% change in profit = (1702500 - 1050000) / 1050000 = 62.14%

Option C will be chosen as it results in the highest increase in profit